OpenStudy (anonymous):

Hey! for 9x^2+y^2-90x+2y+190,where does the derivative =0 and where is it undefined?

OpenStudy (anonymous):

The derivative is not zero since this is not a constant

OpenStudy (anonymous):

oops i wasnt paying attention as usual

OpenStudy (anonymous):

you want \[\frac{dy}{dx}\] yes?

OpenStudy (anonymous):

ready? here we go

OpenStudy (anonymous):

\[18x+2yy'-90+2y'=0\]

OpenStudy (anonymous):

i assume this beast is set = 0, otherwise it does not define a curve, and therefor no implicit function. yes?

OpenStudy (anonymous):

well,its supposed to be a conic equation

OpenStudy (anonymous):

well if it is an "equation" it must be equal something right?

OpenStudy (anonymous):

yeah..it was =to 0

OpenStudy (anonymous):

\[ 9x^2+y^2-90x+2y+190\]as it is is simply an expression.

OpenStudy (anonymous):

how did i guess?

OpenStudy (anonymous):

ok now we solve for y' using elementary algebra

OpenStudy (anonymous):

n its -18x+19/2y+2 right

OpenStudy (anonymous):

i dont get where it wud b undefined?

OpenStudy (anonymous):

supposing it's dy/dx... becomes 0 when x=5

OpenStudy (anonymous):

\[18x+2yy′−90+2y′=0\] \[(2y+2)y'=90-18x\] \[y'=\frac{90-18x}{2y+2}=\frac{45-x}{y+1}\]

OpenStudy (anonymous):

did i make an algebra mistake, because that is what i got

OpenStudy (anonymous):

anyway it is undefined if y = 1

OpenStudy (anonymous):

ohh no u didnt..minz was wrong

OpenStudy (anonymous):

ok then i was just wrong too. it is undefined if y = -1

OpenStudy (anonymous):

ohh..so to find undefined values..make the deniminator = to 0?

OpenStudy (anonymous):

xactly

OpenStudy (anonymous):

i know back to the dungeon for me...

OpenStudy (anonymous):

thnk u soo much!

OpenStudy (anonymous):

welcome. btw 0 if x = 45

OpenStudy (amistre64):

9x^2 +y^2 -90x +2y +190 18x +2y.y' -90 +2.y' y'(2y +2) = -18x +90 \[y' = \frac{-9x +45}{y+1}\] is what I get :)

OpenStudy (anonymous):

ya its undefined when x=-1

OpenStudy (anonymous):

ohhok..thnk u all so much..super appreciate it:)

OpenStudy (amistre64):

\[\frac{dy}{dx}=\frac{-9(x-5)}{y+1}\] y' = 0 when x = 5 and is undefined at x=-1

OpenStudy (anonymous):

it =0 at 90/18 right if the derivative is -18x+90/2y+2

OpenStudy (amistre64):

yes; which simplifies all the way down to 5

OpenStudy (anonymous):

ohyea..oops..ur right..thnks a lottt

OpenStudy (amistre64):

:) your welcome

OpenStudy (anonymous):

Undefined at y=-1 and zero when x=5. You guys said that wrong like 4 times lol.